An unable stunk i mass 1.06 x 18th kg Initially it
Answer #1

M = 1.66*10-26

m1 = 5.20*10-27

m2 = 8.44*10-27

v1 = 6*106 j

v2 = 4*106 i

m3 = M - (m1+m2) = 2.96*10-27

Imomentum = finalmomentum

0 = m1*v1 + m2*v2 + m3*v3

v3 = -(m1*v1 +m2*v2)/m3

v3 = - (11.40*106 i ) - (10.5*106 j )

For part B

initial KE = o
final KE = 0.5[m1v1^2+m2v2^2+m3v3^2]
Solve for the values